Pixel:

what is 2+2+2+3 x 45 / 3 + 65

Eiwoh2:

Final result is 116.

Pixel:

its 200

Pixel:

nub

katelyn66:

i got 200

Eiwoh2:

Lol calculator got 116 XD

AngeI:

116

AngeI:

that was very easy

AngeI:

WARN ME MODS

Eiwoh2:

Oof. When you do it step by step it is 200 lol

Pixel:

angel its not 116

Pixel:

its 200

katelyn66:

you have to do 2+2=4+2=6+3=9 then 9*45=405/3=135+65=200

Eiwoh2:

Yes, but who shall we say is smarter: The Google calculator or a human XD

AngeI:

Kate that is wrong

ThisGirlPretty:

Is this a real math question?

Eiwoh2:

Looks like i have to make some calls to the company. brb.

AngeI:

you have to multiply and divide 1st

Eiwoh2:

Yes, Pretty. It is.

Pixel:

angel

katelyn66:

@angel i dont think so

Pixel:

i want it from left to right

Pixel:

not pemdas

AngeI:

Pixel you cant do it like that or it is wrong x'D

Pixel:

IDC

ThisGirlPretty:

No it isn't.he just did that as a random math question. Please no spam posts.

katelyn66:

@Pixel is it right ?

Eiwoh2:

116 is what Google says,

AngeI:

1 attachment
Eiwoh2:

I, as a person who hates math, would go with that XD

Pixel:

I got it right with 200

Eiwoh2:

Oof. Yep. Calling the company.

Eiwoh2:

Excuse me, Google, but your calculator is not trustworthy. Yeah, i know. Too bad. Fix it, or you will get less visitors XD

AngeI:

Pemdas or its wrong though :/

Eiwoh2:

All in all, this is a trick math question. One of the MANY reasons why i hate mathematics.

Pixel:

i didnt do pemdas and i got it right

AngeI:

-.-

ThisGirlPretty:

@Ultrilliam

katelyn66:

yes, the answer would be 116 if you wanna do it with pemdas but you have to type it in just as you see

Eiwoh2:

Pretty, this is a real math question. We are discussing it. There isn't anything wrong with it, so there is no need to call the teacher.